Intercontinental, Inc., provides you with the following data for its single product Sales price per unit Fixed costs (per month): 50.00 Selling, general, and administrative (SG&A) Manufacturing overhead 900,000 1,800,000 Variable costs (per unit) Direct labor Direct materials Manufacturing overhead 7.00 11.00 10.00 4.00 SG&A Number of units produced per month 300,000 units Required: Compute the amounts for each of the following assuming that both production levels are within the relevant range. (Do not round intermediate calculations. Round your answers to 2 decimal places9 300,000 units 400,000 units a. Prime cost per unit b. Contribution margin per unit C s margin per unit. d. Conversion cost per unit e. Variable cost per unit f. Full absorption cost per unit g. Variable production cost per unit h. Full cost per unit.
Answer #1
Particular 300,000 units
Prime cost per unit $18
Contribution margin per unit $18
Gross margin per unit $16
Conversion cost per unit $23
Variable cost per unit $32
Full absorption cost per unit $34
Variable production cost per unit $28
Full cost per unit $41

Working notes

Prime cost per unit(direct labor + direct material) = $7 + $11 = $18

Contribution margin per unit ( sales Price - variable cost per unit) = $50 - $32 = $18

Variable cost per unit (direct labor + direct material + Manu. Overhead + SG&A) = $7 + $11 + $10 + $4 = $32

Gross margin per unit (selling price - absorption cost) = $50 - $34 = $16

Conversion cost per unit (labor + overhead) = $7 + $10 + ($1,800,000/300,000) = $23

Full absorption cost per unit (Direct labor + direct material +overhead) = $7 + $11 + $10 + ($1,800,000/300,000) = $34

Variable production cost per unit (Direct labor + direct material + Manu. Overhead) = $7 +$11 + $10 = $28

Full cost per unit (Variable cost per unit + fixed cost per unit) = $32 + $9 = $41

Particular 400,000 units
Prime cost per unit $18
Contribution margin per unit $18
Gross margin per unit $17.5
Conversion cost per unit $21.5
Variable cost per unit $32
Full absorption cost per unit $32.5
Variable production cost per unit $28
Full cost per unit $38.75

Working notes

Prime cost per unit(direct labor + direct material) = $7 + $11 = $18

Contribution margin per unit ( sales Price - variable cost per unit) = $50 - $32 = $18

Variable cost per unit (direct labor + direct material + Manu. Overhead + SG&A) = $7 + $11 + $10 + $4 = $32

Gross margin per unit (selling price - absorption cost) = $50 - $32.5 = $17.5

Conversion cost per unit (labor + overhead) = $7 + $10 + ($1,800,000/400,000) = $21.5

Full absorption cost per unit (Direct labor + direct material +overhead) = $7 + $11 + $10 + ($1,800,000/400,000) = $32.5

Variable production cost per unit (Direct labor + direct material + Manu. Overhead) = $7 +$11 + $10 = $28

Full cost per unit (Variable cost per unit + fixed cost per unit) = $32 + $6.75= $38.75